OpenStudy (anonymous):

Choose the correct simplification of the expression a2 • a3. a a6 a5 a−1

OpenStudy (saifoo.khan):

\[\Large a^x * a^y \to a^{x+y}\]

Parth (parthkohli):

What is\[(a \times a) \times (a \times a\times a) \]?

Parth (parthkohli):

Hint: We have a multiplying to itself five times.